Calculate % yield of the reaction if 200g KHCO3 produces 22g of Co2 upon strong heating.

Grade:12

3 Answers

Naveen Kumar
9 years ago
2KHCO3.........>...........K2CO3.+...........CO2.+........H2O
200g....................................................22g


According to the reaction, 2 moles ofKHCO3 should give 1 mole ofCO2.
i.e......................................2*100g.​KHCO3.................................1*44gCO2 for 100% yield.
i.e. 200gKHCO3 should produce.................44g of CO2
but 200g.....................................................22g of CO2
so % yield={actual yield/Theoretical yield}*100
...............={22/44}*100=50%
Pranjul mishra
4 years ago
As masses are given so when we will calculate the moles 
Molar mass of khco3 = 100
Moles are =200/100=2 moles
Moles of co2=2 according to given data but,
The reaction is as follows 
2khco3---->k2co3+co2+h20
So according to reaction 2moles khco3 gives 1 mole co2 so % yeild = given weight /actual weight *100=22/44*100=50%
 
Hope it would be helpful ☺️☺️
 
Hope it would be helpful 
tanyabhardwaj
4 years ago
 Hey, I am looking for a dedicated study partner for JEE. I have put my profile and study schedule on StudyPal. All my friends use it to find study buddies and groups for JEE test